OpenStudy (kylav210):

A wooden plate is shaped as a pentagon. The measure of one angle is 20°. The remaining interior angles are obtuse and of equal measure. What is the measure of each obtuse angle in the pentagon? 140° 155° 175° 130°

OpenStudy (anonymous):

let, interior angles be "x" so, according to question 20+x+x+x+x=540 20+4x=540 4x=540-20 4x=520 x=520/4=130